# Trajectory-based Road Autolabeling with Lidar-Camera Fusion in Winter Conditions

### v1

@@ Abstract
Robust road segmentation in all road conditions is required for safe autonomous driving and advanced driver assistance systems. Supervised deep learning methods provide accurate road segmentation in the domain of their training data but cannot be trusted in out-of-distribution scenarios. Including the whole distribution in the trainset is challenging as each sample must be labeled by hand. Trajectory-based self-supervised methods offer a potential solution as they can learn from the traversed route without manual labels. However, existing trajectory-based methods use learning schemes that rely only on the camera or only on the lidar. In this paper, trajectory-based learning is implemented jointly with lidar and camera for increased performance. Our method outperforms recent standalone camera- and lidar-based methods when evaluated with a challenging winter driving dataset including countryside and suburb driving scenes. The source code is available at https://github.com/eerik98/lidar-camera-road-autolabeling.git
Index Terms:
Trajectory-based learning, Road segmentation, Lidar-camera fusion, Winter driving conditions
I Introduction
@@
In this paper, we propose a novel trajectory-based autolabeling method leveraging both camera and lidar (Fig. 1). We are the first to propose lidar-camera fusion for trajectory-based learning. Our method is validated with a diverse winter driving dataset, demonstrating significant improvements over existing self-supervised methods and even surpassing a supervised baseline trained with manually labeled data.
Figure 1: Proposed trajectory-based road autolabeling with lidar-camera fusion. Trajectory points are determined in the scan and image and labels are generated automatically using the trajectory as a reference. The lidar- and camera-based labels are fused to produce the final label.
II Related work
II-A Trajectory-based learning
@@
II-B Lidar-camera fusion for traversability estimation
Lidar-camera fusion is commonly used for traversability estimation, but none of the fusion methods are trajectory-based. Hand-crafted image and lidar features have been combined for traversability estimation in structured [11] and off-road environments [12]. However, hand-crafted image features are not robust in complex environments with varying road appearance and illumination. In [13] a pre-trained deep-learning-based terrain segmenter is used instead, but the terrain segmenter can’t adapt while driving as it is pre-trained. Lidar can also be used to automatically generate traversability labels for camera-based detection [14, 15, 16]. Here lidar is simply used to generate ground truth and is not used for inference. Lidar-camera fusion is also used with supervised deep-learning models to increase the robustness of road detection in adverse conditions where lidar and camera suffer from weather effects [17].
II-C Vision foundation models
@@
II-D Lidar road boundary detection
@@
II-E Research Gap
In this work, we present a novel trajectory-based autolabeling method leveraging lidar-camera fusion. A prediction model is trained only with the proposed autolabels yielding state-of-the-art performance. Our method is the first trajectory-based approach to utilize both lidar and camera. Our camera-based detection leverages the vision foundation model DinoV2. We compute DinoV2 feature similarity with the trajectory, which provides a continuous estimate for the visual similarity. We have already demonstrated the efficacy of this approach in our previous work [25]. Others [9] have used SAM [20], but it directly outputs segmentation masks that tend to be inaccurate in unstructured environments. Our lidar-based detection incorporates principles of lidar boundary detection with a trajectory-based learning approach. Traditional lidar boundary detection techniques rely on tunable parameters and struggle in unstructured environments [22, 24, 23]. To address this issue, our method identifies trajectory points at each scan ring and uses them as reference points when evaluating the points at that ring.
III Methods
In this section, our novel trajectory-based autolabeling method utilizing both lidar and camera is presented (Fig. 6). Our method first finds the trajectory points in the lidar scan (Section III-A). The lidar wheel points are projected to the image frame and all pixels inside the polygon defined by them are considered as trajectory pixels. Using the lidar trajectory points as a reference a lidar-based autolabel $l^{\text{lid}}$ is computed (Section III-B), and using the trajectory pixels as a reference a camera-based autolabel $l^{\text{cam}}$ is computed (Section III-C). Both autolabels are continuous ranging from 0 to 1, with 1 indicating a definite road and 0 indicating a definite background. The fused label is the mean of the camera-based and lidar-based autolabel. Finally, the fused label is refined with Conditional Random Field (CRF) post-processing [26].
III-A Fit trajectory to scan
The objective is to identify lidar points corresponding to the center, the left wheel, and the right wheel of the vehicle at each scan ring, using recorded vehicle poses (Fig. 2). To locate the center point, we identify the lidar point nearest to the recorded poses within each scan ring. Based on the vehicle’s heading, the estimated positions of the left and right wheels are extended outward from the center point and the lidar points closest to these estimated wheel positions are selected.
@@ III-A Fit trajectory to scan
To avoid false detections following filtering steps are applied. The proposed parameter values were found to work well with our 32-channel lidar.
  • Noise filtering: Statistical outlier filtering is applied using 20 neighbors and a standard deviation ratio of 1.
  • Center point filtering: The center point is rejected if the distance between the pose and the closest scan point is more than 1 m, if the distance to the previous center point is less than 1m, or if the height is more than 1 m above the previous center point (Fig. 3).
  • Wheel point filtering: Wheel point is rejected if the distance between the wheel point and trajectory center is more than 2 m or if it is not visible in the image due to occlusion (Fig. 3).
If either of the wheel points or the center points is filtered out that scan ring is discarded as we don’t have all required reference points. The rest of the scan rings are still used.
@@
Figure 3: Top row: false points from the van are rejected based on the center point filtering rule. Bottom row: occluded wheel points are rejected.
III-B Lidar-based autolabeling
@@
III-B1 Height-based autolabeling
@@
III-B2 Gradient-based autolabeling
@@
III-C Camera-based autolabeling
@@
III-D Data
Our method was validated using a dataset collected in Finland in February 2024. Data acquisition took place across two distinct environments: suburban and countryside roads. The data was gathered with a research vehicle equipped with the following sensors:
  • GNSS/INS: Novatel PWRPAK 7DE2
  • Camera: FLIR blackfly S 2448x2048
  • Lidar: Velodyne VLP-32C
Intrinsic parameters and radial distortion coefficients were calibrated with MATLAB’s checkerboard calibration tool. The extrinsic translation between the lidar and camera frames was measured, and rotations were manually refined to optimize alignment in the validation set. Sensor readings were synchronized based on timestamps and frames where windshield wipers were visible were removed. The trajectory was determined based on the GNSS/INS unit estimates.
For training and evaluation, data sequences from each environment were split into train, validation, and test sets, with unique road segments in each split. Balanced data distribution was achieved by sampling every 5 meters in the train set and every 10 meters in the validation and test sets. Each scene included 100 validation and 400 test images, all manually labeled. The training set consisted of approximately 7k samples for the suburban scene and 10k samples for the countryside scene.
III-E Validation
@@
III-E1 OURS
Trained solely with our autolabels generated for the whole train set with parameters: $\sigma_{\text{C}}=0.6$, $\sigma_{\text{H}}=0.1$, $\sigma_{\text{G}}=0.02$. The prediction model is Deeplabv3 [27] with a ResNet50 backbone, trained with Adam optimizer (batch size 32, 60 epochs, learning rate 1e-3).
III-E2 Supervised
Trained on 200 manually labeled validation images via cross-validation, using the same hyperparameters and model as OURS.
III-E3 SAM2[21]
SAM2 generates a set of segmentation mask proposals around a given input point, each associated with a quality score. We prompt SAM2 with a fixed point taken in front of the vehicle (50 pixels from the bottom, center of the image). In our driving scenarios, this point always lies inside the road. From the generated proposals, we choose the mask with the highest quality score.
III-E4 Seo et. al.[8]
Trained with the whole train set. The learning rate is set to 1e-4 and the batch size to 32, with other parameters as in the original publication.
III-E5 Lidar Boundary Detection[23]
Configured to match our hardware specifications with 32 scan rings and 0.2-degree angular resolution. The ground point segmentation distance threshold was set to 2 m and the scan was limited to a 90-degree field of view, with noise filtered out as with our autolabels. Using the full scan reduced performance. Other parameters follow the original publication.
IV Results
We use IoU to evaluate the overall quality of the road predictions. Precision (PRE), Recall (REC), and F1-score (F1) are also reported.
The efficacy of each proposed autolabeling component was evaluated through an ablation study (Table I). We also present example labels generated with different configurations (Fig. 7). Using only the camera-based label resulted in an IoU of 83.0. Adding the gradient-based label increased IoU by +3.9, while including the height-based label provided an additional +2.8 IoU. Incorporating CRF further improved IoU by +0.4, bringing the final autolabeling accuracy to 90.1 IoU.
Our autolabels can be further evaluated by training a prediction model on them. We chose Deeplabv3 [27] for this purpose, referring to the Deeplabv3 trained with our autolabels as OURS. While lidar data is used in autolabeling, our prediction model is fully image-based. We compare our model’s performance against several baselines (Table II): a zero-shot segmentation model (SAM2 [21]), a lidar boundary detection model [23], a trajectory-based traversability model (Seo et. al.[8]), and a supervised model. We also present examples of predictions for each method (Fig. 8).
@@ IV Results
TABLE I: Ablation study of the proposed autolabeling components for our dataset. Intersection over Union (IoU), Precision (PRE), Recall (REC), and F1-score (F1) are reported. Highest value in bold.
@@
Camera-based (C) 83.0  94.4  87.2  90.7 86.1  96.0  89.3  92.5 80.4  93.1  85.4  89.1
Height-based (H) 80.1  81.6  97.8  89.0 74.5  75.2  98.7  85.4 85.7  88.1  97.0  92.3
Gradient-based no thresholding 45.6  98.1  46.0  62.7 51.4  98.0  52.0  67.9 40.7  98.2  41.0  57.8
Gradient-based (G) 83.5  94.8  87.6  91.0 86.3  93.7  91.6  92.7 81.1  95.7  84.1  89.6
H + G 86.9  91.4  94.6  93.0 86.0  88.0  97.4  92.5 87.7  94.7  92.2  93.4
C + H 83.4  85.1  97.7  91.0 79.0  79.7  98.8  88.3 87.7  90.4  96.7  93.4
C + G 86.2  95.9  89.5  92.6 89.2  95.3  93.4  94.3 83.6  96.4  86.3  91.0
C + H + G 89.7  94.3  94.9  94.6 90.7  92.6  97.8  95.1 88.9  95.8  92.5  94.1
C + H + G + CRF 90.1  94.5  95.1  94.8 90.8  92.8  97.7  95.2 89.4  96.0  92.8  94.4
@@
TABLE II: Prediction performance. Intersection over Union (IoU), Precision (PRE), Recall (REC), and F1-score (F1) are reported. Highest value in bold. Tested on Nvidia RTX 3070.
@@
Lidar Boundary Detection 70.0  74.3  92.3  82.3 68.4  72.1  93.1  81.2 71.4  76.4  91.6  83.3 3.8
@@
V Discussion
This paper introduced a novel trajectory-based lidar-camera fusion method for automatic road labeling, outperforming existing methods on our winter driving dataset. Our lidar-based autolabeling effectively handles various road types by comparing height and gradients with known trajectory points. The camera-based autolabel utilizes DinoV2 feature similarity with the trajectory, leveraging pre-trained features without any need for additional training by the user. Our approach requires few parameters while providing excellent performance in challenging driving scenarios. Both the camera and lidar-based autolabels are continuous rather than discrete, offering a more descriptive estimate of the road state, particularly in unclear boundary situations. Continuous labels also allow meaningful fusion of multiple labels.
Our fusion method also provides increased reliability as it includes two independent autolabeling methods. While lidar-based labels are unreliable when there are no height differences or gradients and camera-based labels are unreliable when the road and roadside appearances are similar, these issues rarely occur together. Our method remains reliable as long as either camera or lidar-based autolabeling provides accurate output.
Currently, our lidar-based autolabeling method assumes that the road elevation is lower than its surroundings and only considers upward height changes for detecting roadside areas. In structured environments and winter driving this assumption usually holds, but for off-road adaption also downward elevation changes should be considered. Also, our method relies on a single scan and image, limiting long-distance labeling capabilities. Future work could integrate consecutive detections to extend the range. Additionally, substituting the mechanical lidar with a stereo camera could reduce hardware costs but modifications to the proposed scan ring-based lidar autolabeling approach would be needed.
VI Acknowledgements

### v2

@@ Abstract
Robust road segmentation in all road conditions is required for safe autonomous driving and advanced driver assistance systems. Supervised deep learning methods provide accurate road segmentation in the domain of their training data but cannot be trusted in out-of-distribution scenarios. Including the whole distribution in the trainset is challenging as each sample must be labeled by hand. Trajectory-based self-supervised methods offer a potential solution as they can learn from the traversed route without manual labels. However, existing trajectory-based methods use learning schemes that rely only on the camera or only on the lidar. In this paper, trajectory-based learning is implemented jointly with lidar and camera for increased performance. Our method outperforms recent standalone camera- and lidar-based methods when evaluated with a challenging winter driving dataset including countryside and suburb driving scenes. The source code is available at https://github.com/eerik98/lidar-camera-road-autolabeling.git. The winter driving dataset is made publically available at https://zenodo.org/records/14856338.
1 Introduction
@@
In this paper, we propose a novel trajectory-based autolabeling method leveraging both camera and lidar. We are the first to propose lidar-camera fusion for trajectory-based learning. Our method is validated with a diverse winter driving dataset, demonstrating significant improvements over existing self-supervised methods and even surpassing a supervised baseline trained with manually labeled data. The main contributions are:
  • 1.
    Utilizing lidar and camera jointly for trajectory-based learning. Existing methods rely only on lidar [10] or camera [8, 7, 9, 11].
  • 2.
    Validation in challenging winter driving conditions. Existing works validated in mostly non-snowy roads. The winter driving dataset is made publically available for future development.
  • 3.
    We outperform both lidar and camera-based baselines with a clear margin. Source code is made publically available for future development. To the best of our knowledge, we are the only trajectory-based method with public source code.
2 Related work
2.1 Trajectory-based learning
@@
2.2 Lidar-camera fusion for traversability estimation
Lidar-camera fusion is commonly used for traversability estimation, but none of the fusion methods are trajectory-based. Hand-crafted image and lidar features have been combined for traversability estimation in structured [12] and off-road environments [13]. However, hand-crafted image features are not robust in complex environments with varying road appearance and illumination. In [14] a pre-trained deep-learning-based terrain segmenter is used instead, but the terrain segmenter can’t adapt while driving as it is pre-trained. Lidar can also be used to automatically generate traversability labels for camera-based detection [15, 16, 17, 18]. Here lidar is simply used to generate ground truth and is not used for inference. Lidar-camera fusion is also used with supervised deep-learning models to increase the robustness of road detection in adverse conditions where lidar and camera suffer from weather effects [19].
2.3 Vision foundation models
@@
2.4 Lidar road boundary detection
@@
2.5 Research Gap
In this work, we present the first trajectory-based approach utilizing both lidar and camera. Our camera-based detection leverages the vision foundation model DinoV2 to evaluate the visual similarity with the trajectory as demonstrated in our previous work [11]. Others [9] have used SAM [22], but it directly outputs segmentation masks that tend to be inaccurate in unstructured environments. Our lidar-based detection incorporates principles of lidar boundary detection with a trajectory-based learning approach. Traditional lidar boundary detection techniques rely on tunable parameters and struggle in unstructured environments [24, 26, 25]. To address this issue, our method identifies trajectory points at each scan ring and uses them as reference points when evaluating the points at that ring.
3 Methods
In this section, our novel trajectory-based autolabeling method utilizing both lidar and camera is presented (Fig. 1). Our method first finds the trajectory points in the lidar scan (Section 3.1). The lidar wheel points are projected to the image frame and all pixels inside the polygon defined by them are considered as trajectory pixels. Using the lidar trajectory points as a reference a lidar-based autolabel $l^{\text{lid}}$ is computed (Section 3.2), and using the trajectory pixels as a reference a camera-based autolabel $l^{\text{cam}}$ is computed (Section 3.3). Both autolabels are continuous ranging from 0 to 1, with 1 indicating a definite road and 0 indicating a definite background. The fused label is the mean of the camera-based and lidar-based autolabel. Beyond lidar’s sensing range we fully rely on the camera-based label. Finally, the fused label is refined to discrete form with Conditional Random Field (CRF) post-processing [27].
3.1 Fit trajectory to scan
The objective is to identify lidar points corresponding to the center, the left wheel, and the right wheel of the vehicle at each scan ring, using recorded vehicle poses (Fig. 2). To locate the center point, we identify the lidar point nearest to the recorded poses within each scan ring. Based on the vehicle’s heading and known track width, the estimated positions of the left and right wheels are extended outward from the center point and the lidar points closest to these estimated wheel positions are selected.
@@ 3.1 Fit trajectory to scan
The lidar scan is limited to a 90-degree field of view as we only need to consider points visible in the camera and the following filtering steps are applied to avoid false detections. The distance between the pose and the selected scan point must be less than 1 m, the distance between consecutive center points must be more than 1 m and the elevation difference between consecutive center points must be less than 1 m (Fig. 3). The wheel point distance from the center point must be less than 2 m and it must not be occluded (Fig. 3). Occlusion filtering is automatic. First lidar scan points (x,y,z) are converted to pixel coordinates (u,v). We consider the line of sight to wheel point $p_{wheel}$ blocked by point $p$ if $|{u_{p}-u_{p_{wheel}}}|<10\text{ pixels}$ and $v_{p}<v_{p_{wheel}}$. If either of the wheel points or the center points is filtered out that scan ring is discarded as we don’t have all required reference points. The rest of the scan rings are still used. The parameter values proposed here were found to work well with our 32-channel lidar.
@@
Figure 3: Top row: False center points rejected based on rapid elevation increase caused by the van. Bottom row: occluded wheel points are rejected.
3.2 Lidar-based autolabeling
@@
3.2.1 Height-based autolabeling
@@
3.2.2 Gradient-based autolabeling
@@
3.3 Camera-based autolabeling
@@
3.4 Data
Our method was validated using a dataset collected in Finland in February 2024. Data acquisition took place across two distinct environments: suburban and countryside roads. The data was gathered with a research vehicle equipped with Novatel PWRPAK 7DE2 GNSS/INS unit, FLIR blackfly S 2448x2048 camera and Velodyne VLP-32C top mounted 3D-lidar. Intrinsic parameters and radial distortion coefficients were calibrated with MATLAB’s checkerboard calibration tool. The extrinsic translation between the lidar and camera frames was measured, and rotations were manually refined to optimize alignment in the validation set. Sensor readings were synchronized based on timestamps and frames where windshield wipers were visible were removed. The trajectory was determined based on the GNSS/INS unit estimates.
For training and evaluation, data sequences from each environment were split into train, validation, and test sets, with unique road segments in each split. Balanced data distribution was achieved by sampling every 5 meters in the train set and every 10 meters in the validation and test sets. Each scene included 100 validation and 400 test images, all manually labeled. The training set consisted of approximately 7k unlabeled samples for the suburban scene and 6k unlabeled samples for the countryside scene. The labeled test and validation samples are made publicly available.
3.5 Validation
@@
OURS is trained solely with our autolabels generated for the whole train set with parameters: $\sigma_{\text{C}}=0.6$, $\sigma_{\text{H}}=0.1$, $\sigma_{\text{G}}=0.02$. The prediction model is Deeplabv3 [28] with a ResNet50 backbone, trained with Adam optimizer (batch size 32, 60 epochs, learning rate 1e-3). Supervised is trained on 200 manually labeled validation images via cross-validation, using the same hyperparameters and model as OURS to demonstrate performance achievable by a limited number of manual labels in contrast to a high number of autolabels. SAM2[23] generates a set of segmentation mask proposals around a given input point, each associated with a quality score. We prompt SAM2 with a fixed point taken in front of the vehicle (50 pixels from the bottom, center of the image). In our driving scenarios, this point always lies inside the road. From the generated proposals, we choose the mask with the highest quality score. Seo et. al.[8] is trained with the whole train set. The learning rate is set to 1e-4 and the batch size to 16, with other parameters as in the original publication. Lidar Boundary Detection[25] is configured to match our hardware specifications with 32 scan rings and 0.2-degree angular resolution. The ground point segmentation distance threshold was set to 2 m and the scan was limited to a 90-degree field of view as with our autolabels. Using the full scan reduced performance. Other parameters follow the original publication.
4 Results
We use IoU to evaluate the overall quality of the road predictions. Precision (PRE), Recall (REC), and F1-score (F1) are also reported. The efficacy of each proposed autolabeling component was evaluated through an ablation study (Table 1). We also present example labels generated with different configurations (Fig. 6). Using only the camera-based label resulted in an IoU of 83.5. Adding the gradient-based label increased IoU by +0.8, while including the height-based label provided an additional +6.5 IoU. Incorporating CRF further improved IoU by +0.4, bringing the final autolabeling accuracy to 90.2 IoU.
Our autolabels were further evaluated by training a prediction model on them. We chose Deeplabv3 [28] for this purpose, referring to the Deeplabv3 trained with our autolabels as OURS. While lidar data is used in autolabeling, our prediction model is fully image-based. We compare our model’s performance against several baselines (Table 2): a zero-shot segmentation model (SAM2 [23]), a lidar boundary detection model [25], a trajectory-based traversability model (Seo et. al.[8]), and a supervised model. We also present examples of predictions for each method (Fig. 7).
@@ 4 Results
Table 1: Ablation study of the proposed autolabeling components for our dataset. Intersection over Union (IoU), Precision (PRE), Recall (REC), and F1-score (F1) are reported. Highest value in bold.
@@
Camera-based (C) 83.5  92.9  89.2  91.0 86.5  94.6  91.0  92.8 81.0  91.5  87.7  89.5
Height-based (H) 80.2  81.8  97.7  89.0 74.4  75.1  98.8  85.3 86.1  88.7  96.7  92.5
Gradient-based no thresholding 34.0  97.6  34.2  50.7 39.0  96.5  39.6  56.1 29.6  98.9  29.7  45.7
Gradient-based (G) 80.5  93.5  85.3  89.2 82.8  91.8  89.4  90.6 78.5  95.2  81.8  88.0
H + G 87.0  92.1  94.0  93.0 86.4  88.5  97.3  92.7 87.5  95.5  91.3  93.4
C + H 83.3  84.8  97.8  90.9 78.5  79.1  99.0  88.0 88.0  90.6  96.8  93.6
C + G 84.3  95.1  88.2  91.5 86.6  93.9  91.7  92.8 82.4  96.2  85.1  90.3
C + H + G 89.8  94.5  94.8  94.6 90.3  92.5  97.5  94.9 89.4  96.3  92.5  94.4
C + H + G + CRF 90.2  94.6  95.0  94.8 90.5  92.6  97.5  95.0 89.9  96.5  92.9  94.7
@@
Table 2: Prediction performance. Intersection over Union (IoU), Precision (PRE), Recall (REC), and F1-score (F1) are reported. Highest value in bold. Tested on Nvidia RTX 3070.
@@
Lidar Boundary Detection 70.2  74.5  92.4  82.5 68.8  72.2  93.6  81.5 71.4  76.6  91.4  83.3 3.8
@@
5 Discussion
This paper introduced a novel trajectory-based lidar-camera fusion method for automatic road labeling, outperforming existing methods on our winter driving dataset. The fusion method provides increased reliability by including two independent autolabeling methods. Most common failure cases are presented in Fig 8. In case a) vibration causes noise to the lidar scan which is detected as a high gradient by the lidar-autolabeling. Good camera-based detection decreases the error in the final label. In case b) camera-autolabeling detects road boundary as road due to visual similarity. Good lidar-based detection decreases the error in the final label. In case c) oncoming lane is not detected by the lidar-autolabeling due to the upward gradients caused by the snowbank. The gradient-based autolabeling can’t handle two separate road segments. In case d) road outside lidar’s range is poorly detected. We rely on a single scan and image limiting long-distance labeling capabilities. In future work consecutive detections could be accumulated for an extended range.
Currently, our lidar-based autolabeling method assumes that the road elevation is lower than its surroundings and only considers upward height changes for detecting roadside areas. In structured environments and winter driving this assumption usually holds, but for off-road adaption also downward elevation changes should be considered. We also assume that the lidar scans aren’t badly degraded. Lidar scan quality can be compromised for example by water layer on top of asphalt. Such scenarios were not included in the study.
Figure 8: Failure cases of the proposed autolabeling method.
6 Acknowledgements